OpenStudy (anonymous):

Simplify 2a+1/3 = a+5

OpenStudy (anonymous):

a = 14/3

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

thats solving not simplifying

OpenStudy (anonymous):

move the a's to the left side and move the fraction to the right side.

OpenStudy (anonymous):

thats not simplifying

OpenStudy (anonymous):

Steps please ! can you draw it or something?

OpenStudy (anonymous):

2a+a=3a 1/3+5=5 3a+5=8a

OpenStudy (anonymous):

i even typed it out

OpenStudy (anonymous):

Thanks